# The simpliciality of higher-order networks

### v2

@@ 2.1 Measures
Formally, we let $\mathbf{H}=(V,E)$ be a hypergraph and let $S=\{e\in E\ |\ \mathcal{P}(e)\subseteq E\}$ be the set of hyperedges which are also simplices. Then, the simplicial fraction is defined as
@@
$\sigma_{\mathrm{ES}}=\frac{|E|}{|C|},$ (2)
again satisfying $\sigma_{\mathrm{ES}}\in[0,1]$; see Fig. 1C. (We note that one can use the induced simplicial complex to define variants of the ES, e.g., a simplicial edit distance $d_{\mathrm{ES}}=|C|-|E|$ or a normalized distance $d_{\mathrm{NES}}=(|C|-|E|)/|C|=1-|E|/|C|=1-\sigma_{\mathrm{ES}}$.)
@@
$\bar{d}_{\mathrm{NFES}}=\frac{1}{\left|F\right|}\sum_{e\in F}\frac{d_{\mathrm{FES}}(e)}{|\mathcal{P}(e)|},$
@@
$\sigma_{\mathrm{FES}}=\frac{1}{\left|F\right|}\sum_{e\in F}\left(1-\frac{d_{\mathrm{FES}}(e)}{\mathcal{P}(e)}\right).$ (3)
We call this last measure the face edit simpliciality (FES).
@@ 3 Results
contact-primary-school 242 12,704 52.50 2.42 0.85 0.92 0.94
contact-high-school 327 7,818 23.91 2.33 0.81 0.93 0.92
hospital-lyon 75 1,824 24.32 2.43 0.91 0.95 0.97
@@
email-enron 143 1,442 10.08 2.97 0.31 0.05 0.50
email-eu 967 23,729 24.54 3.12 0.32 0.05 0.52
@@
congress-bills 1,715 58,788 34.28 4.95 0.03 0.01 0.10
tags-ask-ubuntu 3,021 145,053 48.01 3.43 0.15 0.25 0.46
@@ 3.3 Local measures of simpliciality
Turning to the spatial distribution of simpliciality shown in the insets, we see that the SF and FES find a region of high simpliciality at the network’s core with regions of low simpliciality on its edges. In fact, these two measures are largely in agreement, with a Pearson correlation coefficient of $\rho=0.84$ between the local SF and FES. (The correlation drops to $\rho=0.69$ when comparing the SF and ES). We also observe several nodes for which simpliciality is undefined. These nodes are only connected via minimal faces in their ego-hypergraphs, and these faces are excluded when calculating both potential and actual simplices.
@@
contact-primary-school 0.15 0.15 0.14
contact-high-school 0.22 0.34 0.24
hospital-lyon -0.02 -0.02 -0.01
@@
ndc-substances 0.57 0.65 0.72
diseasome N/A 0.46 0.75
disgenenet N/A 0.55 0.89
@@
congress-bills 0.78 0.48 0.75
tags-ask-ubuntu -0.03 -0.08 0.04
@@
One should expect local simpliciality to be assortative as any given subface contributes to the simpliciality of all their nodes. Table 2 shows that the situation is a bit more nuanced. For tags-ask-ubuntu, FES is weakly assortative, whereas the other two measures are weakly disassortative. It is particularly striking that despite the hospital-lyon dataset being highly simplicial (as seen in Fig. 2), it is also weakly disassortative.
@@ Appendix B Measuring simpliciality efficiently
$\sigma_{\mathrm{SF}}=0$ $F=\{e\in E\ |\ |e|\in K,|e|\geq m\}$ // Iterate over eligible simplices.
@@
       IsSimplex = true for $e\in\mathcal{P}_{K}(f)$ do
@@
                   IsSimplex = false break
@@
$\sigma_{\mathrm{ES}}=0$ // Construct the set of maximal faces.
$F=\{e\in E\ |\ e\notin f,\ \forall f\in E,|e|\geq m\}$ // D stores the unique missing subfaces.
$D=\emptyset$, is a set of sets // Iterate over all maximal faces.
for $f\in F$ do
@@
$\sigma_{\mathrm{ES}}=|E|/(|E|+|D|)$ return $\sigma_{\mathrm{ES}}$
@@
$\sigma_{\mathrm{ES}}=0$ // Construct the set of maximal faces.
$F=(e\in E\ |\ e\notin f,\ \forall f\in E,|e|\geq m)$ $d=0$ // Iterate over all enumerated maximal faces.
@@
       $f=F_{i}$ // First, calculate the number of missing faces for a given maximal face.
       $\widetilde{d}=|\mathcal{P}_{K}(f)|$ for $e\in\mathcal{P}_{K}(f)$ do
@@
       $d\leftarrow d+\widetilde{d}$ // Calculate the number of redundant missing subfaces counted for the maximal faces already seen. To prevent looping over all previous maximal edges, we iterate only over the previous maximal faces, which are also neighbors of the current maximal face.
       $D=\emptyset$ for $j=\{1\dots i-1\}\cap\{k\ |\ e_{k}\cap f\neq\emptyset\}$ do
@@
$\sigma_{\mathrm{ES}}=|E|/(|E|+d)$ return $\sigma_{\mathrm{ES}}$
@@
$\sigma_{\mathrm{FES}}=0$ // Construct the set of maximal faces.
$F=\{e\in E\ |\ e\notin f,\ \forall f\in E,|e|\geq m\}$ $\sigma_{\mathrm{FES}}=0$ // Iterate over all maximal faces.
@@ Acknowledgements
N.W.L. would like to acknowledge the participants of the ”Workshop on Modelling and Mining Complex Networks as Hypergraphs” at Toronto Metropolitan University and Tim LaRock for helpful conversations. N.W.L. would also like to thank Tzu-Chi Yen for lending his expertise on the biSBM inference.

### v3

@@ 2.1 Measures
Formally, we let $\mathbf{H}=(V,E)$ be a hypergraph and let $S=\{e\in E\mid\mathcal{P}(e)\subseteq E\}$ be the set of hyperedges which are also simplices. Then, the simplicial fraction is defined as
@@
$\sigma_{\mathrm{ES}}=\frac{|E|-|\widetilde{E}|}{|C|-|\widetilde{E}|},$ (2)
again satisfying $\sigma_{\mathrm{ES}}\in[0,1]$; see Fig. 1C. We subtract the number of maximal edges from the total number of edges so that the edit simpliciality is zero in the case of a hypergraph with no subfaces. (We note that one can use the induced simplicial complex to define variants of the ES, e.g., a simplicial edit distance $d_{\mathrm{ES}}=|C|-|E|$ or a normalized distance $d_{\mathrm{NES}}=(|C|-|E|)/(|C|-|\widetilde{E}|)=1-(|E|-|\widetilde{E}|)/(|C|-|\widetilde{E}|)=1-\sigma_{\mathrm{ES}}$.)
@@
$\bar{d}_{\mathrm{NFES}}=\frac{1}{\left|F\right|}\sum_{e\in F}\frac{d_{\mathrm{FES}}(e)}{\left|\mathcal{P}(e)\right|-1},$
@@
$\sigma_{\mathrm{FES}}=\frac{1}{\left|F\right|}\sum_{e\in F}\left(1-\frac{d_{\mathrm{FES}}(e)}{\left|\mathcal{P}(e)\right|-1}\right).$ (3)
We call this last measure the face edit simpliciality (FES). We subtract one in the denominators of both expressions so that when an edge has no subfaces, its normalized face edit distance is one.
@@ 3 Results
contact-primary-school 242 12,704 52.50 2.42 0.85 0.88 0.94
contact-high-school 327 7,818 23.91 2.33 0.81 0.91 0.92
hospital-lyon 75 1,824 24.32 2.43 0.91 0.94 0.97
@@
email-enron 143 1,442 10.08 2.97 0.31 0.04 0.50
email-eu 967 23,729 24.54 3.12 0.32 0.04 0.52
@@
congress-bills 1,715 58,788 34.28 4.95 0.03 0.00 0.10
tags-ask-ubuntu 3,021 145,053 48.01 3.43 0.15 0.11 0.46
@@ 3.3 Local measures of simpliciality
Turning to the spatial distribution of simpliciality shown in the insets, we see that the SF and FES find a region of high simpliciality at the network’s core with regions of low simpliciality on its edges. In fact, these two measures are largely in agreement, with a Pearson correlation coefficient of $\rho=0.84$ between the local SF and FES. (The correlation drops to $\rho=0.6$ when comparing the SF and ES). We also observe several nodes for which simpliciality is undefined. These nodes are only connected via minimal faces in their ego-hypergraphs, and these faces are excluded when calculating both potential and actual simplices.
@@
contact-primary-school 0.15 0.17 0.14
contact-high-school 0.22 0.37 0.24
hospital-lyon -0.02 -0.01 -0.01
@@
ndc-substances 0.56 0.54 0.69
diseasome N/A 0.28 0.68
disgenenet N/A 0.28 0.78
@@
congress-bills 0.78 0.33 0.75
tags-ask-ubuntu -0.03 -0.24 0.04
@@
One should expect local simpliciality to be assortative as any given subface contributes to the simpliciality of all their nodes. Table 2 shows that the situation is a bit more nuanced. For tags-ask-ubuntu, FES is weakly assortative, SF is weakly disassortative, and ES is strongly disassortative. It is particularly striking that despite the hospital-lyon dataset being highly simplicial (as seen in Fig. 2), it is also weakly disassortative.
@@ Appendix B Measuring simpliciality efficiently
$\sigma_{\mathrm{SF}}=0$
$F=\{e\in E\mid|e|\in K,|e|\geq m\}$
// Iterate over eligible simplices.
@@
       IsSimplex = true
       for $e\in\mathcal{P}_{K}(f)$ do
@@
                   IsSimplex = false
                   break
@@
$\sigma_{\mathrm{ES}}=0$
$E=\{e\in E\mid|e|\in K,|e|\geq m\}$
// Construct the set of maximal faces.
$F=\{e\in E\mid e\notin f,\ \forall f\in E\}$
// D stores the unique missing subfaces.
$D=\emptyset$, is a set of sets
// Iterate over all maximal faces.
for $f\in M$ do
@@
$\sigma_{\mathrm{ES}}=(|E|-|\widetilde{E}|)/(|E|-|\widetilde{E}|+|D|)$
return $\sigma_{\mathrm{ES}}$
@@
$\sigma_{\mathrm{ES}}=0$
$E=\{e\in E\mid|e|\in K,|e|\geq m\}$
// Construct the set of maximal faces.
$F=\{e\in E\mid e\notin f,\ \forall f\in E\}$
$d=0$
// Iterate over all enumerated maximal faces.
@@
       $f=F_{i}$
      // First, calculate the number of missing faces for a given maximal face.
       $\widetilde{d}=|\mathcal{P}_{K}(f)|$
       for $e\in\mathcal{P}_{K}(f)$ do
@@
       $d\leftarrow d+\widetilde{d}$
       // Calculate the number of redundant missing subfaces counted for the maximal faces already seen. To prevent looping over all previous maximal edges, we iterate only over the previous maximal faces, which are also neighbors of the current maximal face.
       $D=\emptyset$
       for $j=\{1\dots i-1\}\cap\{k\mid e_{k}\cap f\neq\emptyset\}$ do
@@
$\sigma_{\mathrm{ES}}=(|E|-|\widetilde{E}|)/(|E|-|\widetilde{E}|+d)$
return $\sigma_{\mathrm{ES}}$
@@
$\sigma_{\mathrm{FES}}=0$
$E=\{e\in E\mid|e|\in K,|e|\geq m\}$
// Construct the set of maximal faces.
$F=\{e\in E\mid e\notin f,\ \forall f\in E\}$
$\sigma_{\mathrm{FES}}=0$
// Iterate over all maximal faces.
@@ Acknowledgements
N.W.L. would like to acknowledge the participants of the ”Workshop on Modelling and Mining Complex Networks as Hypergraphs” at Toronto Metropolitan University and Tim LaRock for helpful conversations. N.W.L. would also like to thank Tzu-Chi Yen for lending his expertise on the biSBM inference. We thank François Théberge for pointing out inconsistencies between the code and notation in the earlier versions of this manuscript.